Please help

Use technology to find the P-value for the hypothesis test described below. The claim is that for a smartphone carrier's data speeds at airports, the mean is u = 15.00 Mbps. The sample size is n = 17 and the test statistic is t = -2.481.
Answer #1

Solution :

The null and alternative hypothesis is ,

H0 :   = 15.00

Ha :    15.00

n = 17

degrees of freedom = n - 1 = 17 - 1 = 16

t = -2.481.

P( t < -2.481 )

= 0.0123

This is the two tailed test .

p-value = 2 * P( t < -2.481 )

p-value = 2 * 0.0123

p-value = 0.0246